A pendulum makes 90 vibrations in a minute.

a.) Find the period T in seconds.

b.) What is its frequency?


1
Expert's answer
2021-09-28T11:11:33-0400

"a.\\
T= (a/60_s)/90=\\
=\frac{a}{{60}\times{90}}\\
=\frac{a}{5400}\; s"

b.

"F=\frac{1}{T}\\
=\frac{5400}{a}\;\;Hz"